OpenStudy (anonymous):

Simplify: 5sqrt27 3sqrt9 A. 3^1/3 B. 3^4/5 C. 3 D. 3^6/75 E. none of these

OpenStudy (anonymous):

\[\sqrt[5]{27\sqrt[3]{9}}\]

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

what I'd do is write 27 and 9 as powers of 3

OpenStudy (anonymous):

so set it up like \[27^{3}*9^{3}\] ?

OpenStudy (bibby):

no as powers of 3 so \(27 = 3^3\) \(9 = 3^2\)

OpenStudy (anonymous):

okay so what would i do with them?

OpenStudy (anonymous):

plug them into the sqrt instead?

OpenStudy (bibby):

\(\huge \sqrt[5]{27\sqrt[3]{9}}=\sqrt[5]{3^3\sqrt[3]{3^2}}\) you can then use properties of exponents to simplify

OpenStudy (anonymous):

okay im not sure how to do that. can you walk me through it?

OpenStudy (bibby):

yeah. \(\huge \sqrt[a]{b}=b^{\frac{1}{a}}\) so: \(\large \sqrt[3]{3^2}=3^{\frac{2}{3}}\)

OpenStudy (bibby):

apply that to the root: \(\huge \sqrt[5]{3^3\sqrt[3]{3^2}}=\sqrt[5]{3^33^{\frac{2}{3}}}\)

OpenStudy (bibby):

recall that \(a^m*a^n=a^{m+n}\)

OpenStudy (bibby):

so \(\huge 3^3*3^{\frac{2}{3}}=3^{3\frac{2}{3}}\)

OpenStudy (bibby):

we rewrite it as an i mproper fraction and do one last simplification brb piss

OpenStudy (bibby):

so what we have now is \(\huge \sqrt[5]{3^{\frac{11}{3}}}\)

OpenStudy (anonymous):

okay.

OpenStudy (bibby):

use the properties of roots then

OpenStudy (anonymous):

i'm getting none of the answer choices...so E?

OpenStudy (bibby):

yep

OpenStudy (anonymous):

alright, thank you :)